# Beyond Scaling: Measuring and Predicting the Upper Bound of Knowledge Retention in Language Model Pre-Training

### v4

@@ Appendix A Licenses and Terms of Use
Table 3: Twenty templates and their corresponding accuracy values for the $\mathbf{BORN}$ relation on the ours-13b model. [S] represents the subject in the knowledge triple.
Template ACC
[S]’s birth took place in 0.186
[S] was brought into existence in 0.187
[S] was born into the world in 0.190
[S]’s nativity is 0.197
The place in which [S] was given birth to is 0.201
The hometown of [S] is 0.203
The birthplace of [S] is 0.212
[S] was delivered in 0.215
[S]’s cradle was in 0.219
[S]’s entry into the world happened in 0.225
[S] first saw the light in 0.228
[S]’s roots are in 0.230
[S]’s life began in 0.230
[S] was born and raised in 0.237
[S]’s origin is 0.238
[S] entered life in 0.251
[S] came into the world in 0.255
[S] was born in 0.255
[S]’s birth occurred in 0.263
[S]’s starting point was 0.282
Table 4: Several cases with significant deviations from the co-occurrence metric are shown here. $N_{s}$ is the number of paragraphs in which the subject appears, $N_{o}$ is the number of paragraphs in which the object appears, and $N_{s,o}$ is the number of paragraphs in which they co-occur. The first two cases exhibit rare co-occurrence but demonstrate high accuracy, whereas the latter two cases show frequent co-occurrence but exhibit low accuracy.
Subject Question Object $N_{s}$ $N_{o}$ $N_{s,o}$ $ACC$
Rare Co-occurrence but High Accuracy
Csepel SC The headquarters of Csepel SC is in Budapest 398 748,860 172 0.655
Liis Lemsalu Liis Lemsalu is a citizen of Estonia 307 967,207 160 0.570
Frequent Co-occurrence but Low Accuracy
Paris Paris is located in Europe 20,902,571 59,408,155 3,761,610 0.043
Ohio The capital of Ohio is Columbus 11,894,303 3,718,483 1,210,957 0.193
Appendix B Templates for Knowledge Triples
[…]

### v5

@@ Appendix C Case Study
Appendix D Templates for Knowledge Triples
[…]
{ "plays": [ "[S] plays", "[S] performs", "[S] engages in playing", "[S] is an musician of", "[S] is involved in playing", "[S] is a performer of", "[S] executes", "[S] is a practitioner of", "[S] is an player of", "[S] is an artist of", "The genre that [S] performs is", "[S] captivates audiences by performing", "[S] performs the musical style of", "[S] is known for his performances in", "[S]’s musical expertise lies in", "[S] is recognized for playing", "[S] is a notable performer of", "[S]’s artistry is showcased in", "[S]’s musical inclinations are towards", "The musical category that [S] excels in is" ], "occupation": [ "The occupation of [S] is", "[S] works as an", "[S] is employed in the profession of an", "[S] pursues a profession as an", "[S] is a professional", "[S] is engaged in the profession of an", "[S]’s vocation is an", "[S] is an professional", "[S]’s role in life is that of an", "[S]’s professional identity is that of an", "The job of [S] is", "[S] is employed as an", "[S] has a career as an", "The profession of [S] is", "[S]’s line of work is", "[S] makes her career as an", "The career of [S] is", "[S] has chosen to be an", "[S] earns a living as an", "[S]’s professional life revolves around being an" ], "work": [ "[S] works in the field of", "[S] specializes in the field of", "[S]’s work focuses on", "[S] is engaged in the area of", "[S]’s professional activity lies in", "[S]’s expertise is in", "[S] operates within the field of", "[S]’s research is concentrated in", "[S] works extensively in", "[S] has dedicated his career to", "[S] is involved in the study of", "[S]’s field of study is", "[S] contributes to the discipline of", "[S] practices in the area of", "[S]’s work domain is", "[S] conducts his work in", "[S]’s field of expertise is", "[S]’s career revolves around", "[S] works professionally in", "[S]’s occupation relates to" ], "is-located": [ "[S] is located in", "The fatherland of [S] is in", "[S] can be found in", "[S] lives in", "[S] is situated within", "[S] is based in", "[S] resides in", "The country in which [S] lives is", "[S]’s location is in", "[S] stays in", "[S]’s address is in", "[S] is present in", "[S] spends most of his time in", "[S] belongs to", "[S]’s position is in", "You can find [S] in", "[S] is positioned in", "[S] exists in", "[S] stays primarily in", "The country in which [S] is located is" ], "locate": [ "[S] is located in", "The place in which [S] is held is", "[S] is held in", "The place where [S] takes place is", "[S] takes place in", "[S] is hosted in", "The venue for [S] is", "[S] happens in", "[S] is organized in", "[S] is staged in", "[S] occurs in", "The place for [S] is", "The location of [S] is", "[S] is based in", "[S] is set in", "[S]’s venue is in", "[S] is arranged in", "[S] is situated in", "People can take part in [S] in", "[S]’s setting is in" ], "found": [ "[S] was founded in", "[S] originated in", "[S] was established in", "[S] came into existence in", "[S] was formed in", "[S] had its beginnings in", "[S] started in", "[S] was created in", "[S] was initiated in", "[S] took shape in", "[S] was launched in", "[S] began in", "[S] emerged in", "The foundation of [S] took place in", "[S] originated from", "[S] was set up in", "[S] was first formed in", "[S] came to be in", "[S] had its origin in", "[S]’s formation occurred in" ], "born": [ "[S] was born in", "[S] came into the world in", "The birthplace of [S] is", "[S] entered life in", "[S] was delivered in", "[S]’s origin is", "[S] first saw the light in", "[S]’s birth occurred in", "[S] was brought into existence in", "[S]’s nativity is", "The place in which [S] was given birth to is", "[S] was born and raised in", "[S]’s roots are in", "The hometown of [S] is", "[S]’s cradle was in", "[S]’s entry into the world happened in", "[S]’s birth took place in", "[S] was born into the world in", "[S]’s life began in", "[S]’s starting point was" ], "die": [ "[S] died in", "[S] passed away in", "[S]’s death occurred in", "[S] lost his life in", "[S]’s demise happened in", "[S] perished in", "[S]’s end came in", "[S]’s life ended in", "[S] met his end in", "[S]’s passing was in", "[S]’s death took place in", "[S]’s final moments were in", "[S]’s last breath was in", "[S]’s life was lost in", "[S]’s expiration occurred in", "[S]’s fatality was in", "[S]’s death scene was in", "The place in which [S] passed away is", "[S]’s death location is", "[S]’s death site is" ], "air": [ "[S] was originally aired on", "[S] first aired on", "[S] premiered on", "[S] was initially broadcast on", "[S] debuted on", "[S] was first shown on", "[S] was originally broadcast on", "[S] was first transmitted on", "[S] was first telecast on", "[S] was first screened on", "[S] was first presented on", "[S] had its first airing on", "[S] was first televised on", "[S] was first released on", "[S] was first featured on", "[S] was first displayed on", "[S] was first exhibited on", "[S] was first put on air on", "[S] was first launched on", "[S] was first introduced on" ], "headquarter": [ "The headquarters of [S] is in", "[S]’s headquarters is in", "[S]’s main office is in", "[S]’s central office is in", "[S]’s head office is in", "The corporate office of [S] is in", "[S]’s principal office is in", "[S]’s administrative center is in", "[S]’s base of operations is in", "[S]’s HQ is in", "The headquarters’s location of [S] is in", "[S]’s headquarters office is in", "[S]’s primary office is in", "[S]’s main administrative office is in", "[S]’s central headquarters is in", "[S]’s main base is in", "The main hub of [S] is in", "[S]’s main building is in", "[S]’s headquarters is located in", "The place in which [S]’s headquarters is situated in" ], "capital": [ "[S] is the capital of", "[S] serves as the capital of", "[S] is recognized as the capital of", "[S] stands as the capital of", "[S] is known as the capital of", "[S] functions as the capital of", "[S] is officially the capital of", "[S] holds the title of capital of", "[S] is designated as the capital of", "[S] is the administrative capital of", "[S] is the political capital of", "[S] is the governmental capital of", "[S] is the principal capital of", "[S] is the main capital of", "[S] is the chief capital of", "[S] is the primary capital of", "[S] is the foremost capital of", "The country whose capital is [S] is", "[S] is known to the people as the capital of", "As the capital city, [S] has an extremely important place in" ], "citizen": [ "[S] is a citizen of", "The country where [S] holds citizenship and enjoys benefits is", "The nation that grants [S] the rights and privileges of citizenship is", "[S] is from the country of", "The place of [S]’s birth and upbringing is", "[S] belongs to", "[S]’s homeland is", "[S]’s home country is", "[S] is a resident of", "[S]’s origin is", "The land where [S]’s roots run deep is", "[S]’s place of origin is", "[S]’s country of origin is", "The nation that [S] calls home is", "[S]’s native country is", "[S]’s nationality is from", "The place where [S] was born and raised is", "[S]’s citizenship gives him the rights and privileges of a citizen in", "[S]’s citizenship makes him a part of the country’s community in", "The country that [S] resides in and holds citizenship is" ], "create": [ "[S] was created in", "The birthplace of [S] is", "[S] originated in", "The invention of [S] took place in", "[S] hails from", "The country that [S] was made in is", "[S] was initially developed in", "The country that introduced [S] is", "[S] was crafted in", "The place of origin for [S] is", "[S] was first produced in", "The homeland of [S] is", "The source of [S] is", "[S] was manufactured in", "The origin of [S]’s creation is", "[S] was designed in", "The place where [S] was conceived is", "[S] was made in", "The birth nation of [S] is", "[S] was developed in" ], "play": [ "[S] plays", "[S] is known for playing the", "[S]’s instrument is the", "[S] is a musician who plays the", "[S] is one who has mastered the art of the", "[S] has a gift for playing the", "Few can rival [S]’s prowess on the", "[S] is a true virtuoso of the", "[S]’s music is infused with the soulful sound of the", "Audiences are captivated by [S]’s", "[S]’s musical talents shine through his", "[S] is a skilled", "One can hear [S]’s passion in every", "With every note, [S] demonstrates his mastery of the", "[S] is a", "[S] plays the", "[S] specializes in the", "[S] performs on the", "Music flows through [S] and his", "As for instruments, [S] prefers the" ], "is-locate": [ "[S] is located in", "[S] lies in the", "[S] lies within the boundaries of", "The place where you can find [S] is", "[S] is situated in", "[S] is positioned in", "[S] exists in", "[S] is placed in", "[S] resides in", "The place where [S] can be fully appreciated is", "[S] can be found in", "[S] is in the area of", "[S] is part of the landscape of", "[S] is in the lands of", "The region that holds [S] is", "The location of [S] is", "[S] is in the sector of", "The place that [S] is a part of is", "[S] is in the section of", "[S] is in the region of" ] }